The heat energy needed to raise the temperature of steam from 100 to 120 degrees Celcius (in Joules) =


1
Expert's answer
2022-01-17T09:58:58-0500

The specific heat capacity of steam is 1.996 J/goC.


The mass of steam is not given, hence it is impossible to solve this problem. The heat depends on the mass of a substance.


For 1 gram of the steam, the heat energy will be following:

"Q=cm\Delta{T}=1.996\frac{J}{g\cdot\degree{C}}\times1\ g\times(120\degree{C}-100\degree{C})=40\ J"